1. AXY Company is expected to pay an annual dividend of $1.6 per share next year and then pay a final liquidating divided of $25 per share the following year. Currently, you own 350 shares of the Golden Cage. What is the current value of your shares if the required return is 10 percent?
Answer #1

Current price = Presnet value of future cash flows

Current price = 1.6 / (1 + 0.1)1 + 25 / (1+ 0.1)2

Current price = 1.454545 + 20.661157

Current price = 22.1157

Current value = 350 * 22.1157

Current value = $7,740.50